Tickets of a program at college cost $3 for general admission or $2 with a student ID. If 179 people paid to see a performance and $431 was colleceted, how many of each type of ticket were sold?
Answer #1

Let x number of people paid $3 and y number of people paid $2. Total $431 is collected.

So, 3x + 2y = 431 ------------- equation-1

Total number of people is 179.

So, x+ y = 179 ------------- equation-2

Solving both equation,

multiplying equation-2 by 2, 2x + 2y = 358 and subtracting from equation-1

(3x + 2y) - (2x + 2y) = 431 -358

=> x = 73,

x = 73 and y = 106

Hence, 73 tickets are sold of type $3 and 106 tickets are sold of $2.